OpenStudy (anonymous):

The surface area of a rectangular prism is 442 square feet. The length is 11 feet, and the height is 8 feet. What is the width of the prism? & Show how you got it.

OpenStudy (campbell_st):

so 442/2 = 221 since 3 pairs of identical faces. Let w be the missing length then 221 = 8 x 11 + w x 11 + w x 8 221 = 88 + 19 w solve the equation to find the width

OpenStudy (anonymous):

Thank you :D